Determine the area to the left of g(y) = 3-y2 and to the right of x = -1

1
Expert's answer
2022-05-22T17:40:35-0400
"-1=3-y^2"

"y_1=-2, y_2=2"

"A=\displaystyle\int_{-2}^{2}(3-y^2-(-1))dy=[4y-\dfrac{y^3}{3}]\begin{matrix}
 2 \\
 -2
\end{matrix}"

"=8-\dfrac{2^3}{3}-(-8+\dfrac{2^3}{3})=\dfrac{32}{3}({units}^2)"